# Lower-mass-gap Black Holes in Dense Star Clusters

### v1

@@ 5 Lower-mass-gap Binaries
Having established that LMG BHs can form in GCs through various mechanisms, we now explore the subsequent dynamical evolution of these objects, in particular the dynamical formation of LMG binaries. In general, only a very small number of NSāNS and BHāNS binaries are present at late times for all cluster simulations, as shown in TableĀ 3.2 (up to about 10 in the extremely optimistic simulationĀ 14). This is consistent with the fact that, despite the hundreds of MSPs observed in Milky Way GCs, only two pulsar-NS (or low-mass BH) binaries have been definitively identified. Both are in massive, core-collapsed, or near core-collapsed clusters (M15 and NGC 1851). We list the properties of the MSP NGCĀ 1851E and its binary in TableĀ 3, together with those of the confirmed or candidate NSāNS binaries in all Milky Way GCs for reference.
@@
On the other hand, the component mass distribution of the NSāNS and BHāNS binaries differs between the two simulations with different mass scenarios for WD mergers and supernova prescriptions. FigureĀ 5 shows the primary and secondary masses of all NSāNS and NSāBH binaries between 10 and 13Ā Gyr of the clustersā evolution from both NGC 1851-like models. The component masses of the NGCĀ 1851E binary are plotted for comparison. Simulation A produces no BHāNS binaries with primary mass in the LMG. In contrast, for () of the time steps between and Ā Gyr, the NSāNS or BHāNS binaries in simulation B have primary mass in the range (), including binaries with masses similar to the NGCĀ 1851E binary (bottom panel of FigureĀ 5). Furthermore, simulation B also produces more massive NSs that can better match the mass of NGCĀ 1851E. This figure suggests that the delayed supernova prescription (or, more generally, the formation of LMG objects from massive star evolution and their retention in dense star clusters) is important for producing LMG binaries in dense and massive clusters similar to NGCĀ 1851. This is because massive clusters can retain more LMG BHs due to their larger escape velocities (TableĀ 4). Furthermore, it also indicates that conserving mass during WDāWD mergers (the updated mass scenario in SectionĀ 2) is a key factor for clusters to produce NGC 1851E-like binaries containing an LMG object and a massive NS companion (see also, e.g., Ye etĀ al., 2024, their FigureĀ 9).
Many NS/BHāNS binaries between 10 and 13Ā Gyr of simulation B have component masses and orbital properties similar to the NGCĀ 1851E binary. In particular, for binaries with primary masses in the range of and secondary masses in the range of (approximately the same as the components in the NGCĀ 1851E binary), the mean and median eccentricities are about 0.7, and the mean and median orbital periods are about 31 and 13 days, respectively 333Note that these values come from only one simulation. It is possible that the values could change slightly with more simulations using the same initial conditions but different random seeds..
@@ 6.1 Gravitational Wave Sources
In general, only a small fraction of binary BH mergers from GCs (a few percent) include LMG BHs between and . However, of merging BHāBH from a typical core-collapsed GC can contain at least one LMG BH if NSs are able to accrete very efficiently from TDEs of main-sequence stars and collapse to BHs (simulation 13 in TableĀ 4). These BHs would be fast-spinning and may be detectable by LVK. On the contrary, of BHāNS mergers involve a BH in the LMG for most simulations in TableĀ 4. The recent detection of GW230529, a merger between a LMG object and an NS from the fourth observing run of LVK, suggests that about of BHāNS mergers may include a component in the LMG (The LIGO Scientific Collaboration etĀ al., 2024). This fraction is similar to what we find for BHāNS mergers in core-collapsed clusters, but the merger rates of GW230529-like systems, (The LIGO Scientific Collaboration etĀ al., 2024), are much larger than the rates expected from GCs. Assuming that about of GCs are core-collapsed (similar to the percentage in the Milky Way) and the GC number density is about (Rodriguez etĀ al., 2016a), the merger rates of all BHāNS mergers from GCs are roughly , even when adopting the delayed supernova prescription and the updated WD mass scenario (the actual rate is probably even lower; Ye etĀ al., 2020, see). The extreme upper limit of the BHāNS merger rates is a factor of a few higher and roughly comparable to the BHāBH merger rates, assuming that all NSs and BHs receive small natal kicks, and that not many BHs with masses greater than are formed in clusters (simulationĀ 14). Therefore, even though we can roughly reproduce the observed MSPāLMG object in NGCĀ 1851, the estimated merger rate of BHāNS is still likely orders of magnitude lower than inferred from the most recent LVK data release (; Abbott etĀ al., 2023).
@@ 7 Discussion
One of the main uncertainties of the formation and evolution of LMG objects in dense star clusters lies in the number of systems retained, which is determined by the supernova mechanism and the kicks low-mass BHs and NSs receive at birth. The escape velocity of typical GCs () imposes an upper limit on the retention fraction, depending on the natal kick distribution. For NS natal kicks estimated from isolated pulsar observations in the Galactic field (e.g., Hobbs etĀ al., 2005, adopted in this study), few massive NSs and low-mass BHs (less than about ) are retained (TableĀ 4), even though are formed in a typical GC assuming the delayed supernova mechanism (Fryer etĀ al., 2012). More recent studies, either based on observed NS binaries (e.g., OāDoherty etĀ al., 2023) or on three-dimensional supernova simulations and semi-analytical models (e.g., Mandel & Müller, 2020), have instead suggested that NSs may receive smaller natal kicks than previously estimated (e.g., FigureĀ 1 in OāDoherty etĀ al., 2023, and references therein). If this is the case, up to times more NSs may be retained in GCs, considering also that most massive stars are born in binaries (Sana etĀ al., 2012; Moe & Di Stefano, 2017). Assuming that LMG BHs, if formed, also receive small natal kicks, an order of magnitude more LMG BHs could be retained in GCs. Together with the larger number of retained NSs, this might lead to a factor of a few more binaries containing LMG BHs in the clusters in the present day. Future (non)observations of LMG binaries will shed more light on the supernova and dynamical formation of these objects in dense star clusters.
@@ 8 Conclusions
In this study, we have explored for the first time the formation and evolution of LMG objects (in the mass range around ) and their binaries through the interplay between massive star evolution and dynamical interactions in dense star clusters. Motivated by the recent detection of a pulsar binary with a potential LMG companion in the massive and dense Milky Way cluster NGCĀ 1851 (Barr etĀ al., 2024), we model this cluster together with typical core-collapsed GCs using our Monte Carlo code CMC, which self-consistently takes into account stellar (and binary star) evolution as well as all -body dynamics.
@@
The delayed supernova prescription produces LMG objects in a typical core-collapsed GC (while the rapid prescription produces none). However, with the large natal kicks they likely receive, only a few LMG BHs are retained in the clusters after formation. More massive GCs similar to NGCĀ 1851 can retain a few tens of these LMG BHs born in delayed supernovae because of the clusterās deeper potential well and the larger number of compact remnants formed initially.

### v2

@@ 4 Formation of Lower-mass-gap Objects
āSimulationĀ C shows a limiting case with Galactocentric distance Ā kpc. Other initial conditions are the same as in simulationĀ B.
@@ 5 Lower-mass-gap Binaries
Having established that LMG BHs can form in GCs through various mechanisms, we now explore the subsequent dynamical evolution of these objects, in particular the dynamical formation of LMG binaries. In general, only a very small number of NSāNS and BHāNS binaries are present at late times for all cluster simulations, as shown in TableĀ 3.2 (average of 6 and up to about 10 in the extremely optimistic simulationĀ 14). This is roughly consistent with the fact that, despite the hundreds of MSPs observed in Milky Way GCs (and found in simulations, e.g., Ye etĀ al., 2020, 2022), only two pulsar-NS (or low-mass BH) binaries have been definitively identified. Both are in massive, core-collapsed, or near core-collapsed clusters (M15 and NGC 1851). We list the properties of the MSP NGCĀ 1851E and its binary in TableĀ 3, together with those of the confirmed or candidate NSāNS binaries in all Milky Way GCs for reference.
@@
On the other hand, the component mass distribution of the NSāNS and BHāNS binaries differs between the two simulations with different mass scenarios for WD mergers and supernova prescriptions. FigureĀ 5 shows the primary and secondary masses of all NSāNS and NSāBH binaries between 10 and 13Ā Gyr of the clustersā evolution from both NGC 1851-like models. The component masses of the NGCĀ 1851E binary are plotted for comparison. Simulation A produces no BHāNS binaries with primary mass in the LMG. In contrast, for () of the time steps between and Ā Gyr, the NSāNS or BHāNS binaries in simulation B have primary mass in the range (), including binaries with masses similar to the NGCĀ 1851E binary (bottom panel of FigureĀ 5). Furthermore, simulation B also produces more massive NSs that can better match the mass of NGCĀ 1851E. This figure suggests that the delayed supernova prescription (or, more generally, the formation of LMG BHs from massive star evolution and their retention in dense star clusters) is important for producing LMG binaries in dense and massive clusters similar to NGCĀ 1851. This is because massive clusters can retain more LMG BHs due to their larger escape velocities (TableĀ 4). Furthermore, it also indicates that conserving mass during WDāWD mergers (the updated mass scenario in SectionĀ 2) is a key factor for clusters to produce NGC 1851E-like binaries containing an LMG BH and a massive NS companion (see also, e.g., Ye etĀ al., 2024, their FigureĀ 9).
Many NS/BHāNS binaries between 10 and 13Ā Gyr of simulation B have component masses and orbital properties similar to the NGCĀ 1851E binary. In particular, for binaries with primary masses in the range of and secondary masses in the range of (approximately the same as the components in the NGCĀ 1851E binary), the mean and median eccentricities are about 0.7, and the mean and median orbital periods are about 31 and 13 days, respectively 555Note that these values come from only one simulation. It is possible that the values could change slightly with more simulations using the same initial conditions but different random seeds..
@@ 6.1 Gravitational Wave Sources
In general, only a small fraction of binary BH mergers from GCs (a few percent) include LMG BHs between and . However, of merging BHāBH from a typical core-collapsed GC can contain at least one LMG BH if NSs are able to accrete very efficiently from TDEs of main-sequence stars and collapse to BHs (simulation 13 in TableĀ 4). These BHs would be fast-spinning and may be detectable by LVK. On the contrary, of BHāNS mergers involve a BH in the LMG for most simulations in TableĀ 4. The recent detection of GW230529, a merger between a LMG BH and an NS from the fourth observing run of LVK, suggests that about of BHāNS mergers may include a component in the LMG (The LIGO Scientific Collaboration etĀ al., 2024). This fraction is similar to what we find for BHāNS mergers in core-collapsed clusters, but the merger rates of GW230529-like systems, (The LIGO Scientific Collaboration etĀ al., 2024), are much larger than the rates expected from GCs. Assuming that about of GCs are core-collapsed (similar to the percentage in the Milky Way) and the GC number density is about (Rodriguez etĀ al., 2016a), the merger rates of all BHāNS mergers from GCs are roughly , even when adopting the delayed supernova prescription and the updated WD mass scenario (the actual rate is probably even lower; Ye etĀ al., 2020, see). The extreme upper limit of the BHāNS merger rates is a factor of a few higher and roughly comparable to the BHāBH merger rates, assuming that all NSs and BHs receive small natal kicks, and that not many BHs with masses greater than are formed in clusters (simulationĀ 14). Therefore, even though we can roughly reproduce the observed MSPāLMG BH in NGCĀ 1851, the estimated merger rate of BHāNS is still likely orders of magnitude lower than inferred from the most recent LVK data release (; Abbott etĀ al., 2023).
@@ 7 Discussion
One of the main uncertainties of the formation and evolution of LMG BHs in dense star clusters lies in the number of systems retained, which is determined by the supernova mechanism and the kicks low-mass BHs and NSs receive at birth. The escape velocity of typical GCs () imposes an upper limit on the retention fraction, depending on the natal kick distribution. For NS natal kicks estimated from isolated pulsar observations in the Galactic field (e.g., Hobbs etĀ al., 2005, adopted in this study), few massive NSs and low-mass BHs (less than about ) are retained (TableĀ 4), even though are formed in a typical GC assuming the delayed supernova mechanism (Fryer etĀ al., 2012). More recent studies, either based on observed NS binaries (e.g., OāDoherty etĀ al., 2023) or on three-dimensional supernova simulations and semi-analytical models (e.g., Mandel & Müller, 2020), have instead suggested that NSs may receive smaller natal kicks than previously estimated (e.g., FigureĀ 1 in OāDoherty etĀ al., 2023, and references therein). If this is the case, up to times more NSs may be retained in GCs, considering also that most massive stars are born in binaries (Sana etĀ al., 2012; Moe & Di Stefano, 2017). Assuming that LMG BHs, if formed, also receive small natal kicks, an order of magnitude more LMG BHs could be retained in GCs. Together with the larger number of retained NSs, this might lead to a factor of a few more binaries containing LMG BHs in the clusters in the present day. Future (non)observations of LMG binaries will shed more light on the supernova and dynamical formation of these objects in dense star clusters.
@@ 8 Conclusions
In this study, we have explored for the first time the formation and evolution of LMG BHs (in the mass range around ) and their binaries through the interplay between massive star evolution and dynamical interactions in dense star clusters. Motivated by the recent detection of a pulsar binary with a potential LMG companion in the massive and dense Milky Way cluster NGCĀ 1851 (Barr etĀ al., 2024), we model this cluster together with typical core-collapsed GCs using our Monte Carlo code CMC, which self-consistently takes into account stellar (and binary star) evolution as well as all -body dynamics.
@@
The delayed supernova prescription produces LMG BHs in a typical core-collapsed GC (while the rapid prescription produces none). However, with the large natal kicks they likely receive, only a few LMG BHs are retained in the clusters after formation. More massive GCs similar to NGCĀ 1851 can retain a few tens of these LMG BHs born in delayed supernovae because of the clusterās deeper potential well and the larger number of compact remnants formed initially.